OpenStudy (anonymous):

Identify the equation in point-slope form for the perpendicular bisector of the segment with end points E(−7, 6) and G(9, −2). I JUST CANNOT REMEMBER HOW TO DO THIS PLEASE HELP

OpenStudy (anonymous):

Well, find the slope of the segment you have currently first. Do you know how to do that?

OpenStudy (anonymous):

@OliviannaC

OpenStudy (anonymous):

yeah I think so, so you take (-2)-6/9-(-7) ?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

I got -11/16...

OpenStudy (anonymous):

Alright, so now you've got the slope of the first equation. You don't have the y-intercept, but you have a point. This means you have to use the point-slope format. Do you remember what it looks like?

OpenStudy (anonymous):

Wot. How did you get that? And no point slope is \[y=m(x-x _{1})+y _{1}\]

OpenStudy (anonymous):

Oh ok, you deleted that post

OpenStudy (anonymous):

which is it? y-2=2(x-1) y-1=2(x-2) y-2=1/2(x-1) y-1=1/2(x-2)

OpenStudy (anonymous):

You sound like you're in hurry so I'll go ahead and give you the answer. One second

OpenStudy (anonymous):

thank you and sorry! I am trying to remember the form. I wrote it down

OpenStudy (anonymous):

Hey if I don't respond in bit it's because it's stormy here and my internet is sketchy

OpenStudy (anonymous):

you said it was y=m(x-x1) y1 but the first x is x2 i am assuming?

OpenStudy (anonymous):

and you can just answer A, B, C, or D of the four answer options

OpenStudy (anonymous):

No, just x, if you multiply it out, it'll be whatever x times the slope mx

OpenStudy (anonymous):

ok ok, that makes sense

OpenStudy (anonymous):

y-2=2(x-1) y-1=2(x-2) y-2=1/2(x-1) y-1=1/2(x-2

OpenStudy (anonymous):

Wait a bit, I'm kind of in the middle of some other things. How soon do you need this?

OpenStudy (anonymous):

i need to get off the computer right now and pack for leaving another country tomorrow...so kinda fast!! :-/

OpenStudy (anonymous):

idk i think i am just having a brain fart right now cause i already know how to do this...

OpenStudy (anonymous):

I'll tag some people for you because I've got something on the stove so I really can't do this right away @AravindG @agent0smith @AccessDenied @sourwing

OpenStudy (anonymous):

ok thanks

OpenStudy (accessdenied):

Your slope of the line was: m = (-2 - 6) / (9 - -7) = -8/16 = -1/2. Because we want a perpendicular slope, we take the opposite reciprocal: m= - 1 / (-1/2) = 2 Then we must find a point. Because it is the perpendicular bisector, we require a midpoint. M ((x1 + x2)/2, (y1 + y2)/2) for the two points. M((9 + -7)/2, (-2 + 6)/2) Simplify for the midpoint, which is where our perpendicular bisector must cross the line. And finally, plug in the information for point-slope form: y - y1 = m (x - x1)